# iFANnpp: Nuclear Power Plant Digital Twin for Robots and Autonomous Intelligence

### v2

@@ Abstract
Robotics has gained significant attention in the nuclear industry due its precision and ability to automate tasks. However, the increasing complexity of robots has led to a growing demand for advanced simulation and control methods to predict robot behavior and optimize plant performance, motivating the use of digital twins for robotic applications. Most existing digital twins only address parts of systems and do not offer a total design of a nuclear power plant. Furthermore, they are often designed for specific algorithms or tasks, making them unsuitable for broader research applications or projects. In response, this work proposes a comprehensive nuclear power plant digital twin designed to improve real-time monitoring, operational efficiency, and predictive maintenance. The full nuclear power plant is modeled in Unreal Engine 5 to incorporate the complexities and various phenomena. The high-resolution simulation environment is integrated with a Generic Pressurized Water Reactor Simulator, a high-fidelity physics-driven software, to create a realistic model of a nuclear power plant and a real-time updated virtual environment. The virtual environment provides various features for researchers to easily test custom robot algorithms and frameworks, applicable to research in the nuclear industry as well as industrial systems in general. The digital twin’s performance is presented, and critical research problems are addressed, including multi-robot task scheduling and robot navigation in radiation-affected areas, by leveraging implemented features.
@@
Digital Twin, Robot Simulation, Nuclear Power Plant, Virtual Reality
@@ I Introduction
Robots move and operate with significant speed and force, making them challenging to train and control effectively. Errors in robot control software can result in severe hardware damage or even injury to humans. In addition, accounting for environmental factors and unknown variables in the real world makes them difficult to test and eventually deploy. To avoid these issues, digital twin or virtual simulation platforms can be relied on for experiments, including risk mitigation, rapid prototyping, and cost-efficiency.
A digital twin is a solution to develop a realistic, virtual model of a system with continuous updates of its real-time data and behavior. It enables real-time monitoring and physics simulation, allowing operators to predict system behaviors. The advancement in predictability improves nuclear power plant reliability by reducing plant failures. Additionally, digital twins provide a platform for testing research ideas without accessibility constraints. Access to nuclear power plants is highly restricted due to their potential hazards and confidentiality. environments and components. A digital twin - a realistic platform - is perfect for this environment since it presents the simulation of the nuclear power plants and produces data that can be used further after running the simulation. Additional follow-up research can be conducted, including simulation of various operations in nuclear power plants, optimization of the operations, prediction of outcomes using robots, and real-time monitoring of nuclear power plant component behaviors.
Figure 1: iFANnpp Architecture and included features
This paper introduces a nuclear power plant digital twin for robots and autonomous intelligence, called iFANnpp, capable of supporting the development of advanced technology such as artificial intelligence, reinforcement learning (RL) and autonomous robot controls. Figure 1 illustrates the overview framework of iFANnpp, which includes a bidirectional communication pipeline for a high-fidelity physics simulator for accurate industrial simulation and Python-Unreal bridge for a user-friendly algorithm testbed. This platform uses Unreal Engine 5, a powerful 3D modeling and simulation software. It offers several key features:
@@
The remainder of the paper is structured as follows: Section II introduces related research on industrial digital twins and associated robots. Section III discusses the modeling of iFANnpp. Section IV describes the various features presented in iFANnpp. Section V demonstrates test scenarios conducted on the digital twin. Section VI presents an analysis on the performance of the digital twin. Section VII concludes the paper.
@@ II-A Industrial Digital Twin
Several studies have explored integrating advanced 3D models with real-time data to create accurate and functional digital twins. Sørensen et al. [14] developed a digital twin using Unreal Engine 5 specifically for wind power analysis, allowing real-time simulation and turbine performance optimization. Garg et al. [15] implemented a virtual reality (VR) feature with Unity 3D to connect a virtual world to a physical FANUC robot, providing remote control and monitoring its operation. Erdei et al. [16] used Unreal Engine 4 to create a cyber-laboratory space for KUKA KR5 work area for training purposes. Metzner et al. [17] created a VR digital twin of a robot in an industrial control system, specifically for simulating programmable logic controllers (PLCs) in Unity. Gonzalez et al. [18] modeled a laboratory room virtually and augmented it with reality capabilities using a mobile device, enhancing virtual interaction more realistically. Vairagade et al. [19] developed an NVIDIA Omniverse-based digital twin for nuclear power plants using a high-fidelity simulator. While similar to this work in concept, their work was limited to realistically modeling a single building within the nuclear power plant system. In contrast, this work models all primary functional components of a pressurized water reactor (PWR) and offers a broader range of features, including bidirectional communication with the underlying simulator and the Python-Unreal bridge.
@@ II-B Robots in Industrial Environments
While the former works focus on the modeling and control of stationary robots to enable fine-grained manipulation and automation for industrial applications, additional works feature mobile robotics. Mobile robotics offer the flexibility of movement, but can also add design complexity regarding power management and algorithmic control [23]. Naveen et al. [24] designed and implemented a solar-powered mobile robot with pick and place capabilities as well as control software for potential industrial automation tasks. Ghodsian et al. [25] provides a survey of industrial mobile robotics, highlighting the importance of their development for Industry 4.0 designs, their flexibility in terms of movement and coordination, as well as the complexities and challenges in their design to ensure valid function and safety. Buerkle et al. [26] discuses a potential industrial robot system architecture using mobile robots, called Industrial Robots as a Service, highlighting potential methods for programming mobile robots such as through explicit task programming or via human-robot interaction. Additionally, Buierkle et al. emphasize the need for robot adaptability to environmental change and the critical importance of safety, primarily when humans work near autonomous machines.
iFANnpp’s realistically modeled environment provides a platform for simulating stationary and mobile robots for nuclear power plant applications. The transition to deploying mobile robots for industrial applications presents many challenges, including the need for more complex algorithm development and a safe environment. Using iFANnpp, researchers can prototype mobile robot planning and control methods, confirming design and execution validity before deploying them to the real world.
@@ III Digital Twin Modeling
iFANnpp includes 3D models for the entire nuclear power plant, shown in Figure 2. A nuclear power plant is one of the most complex industrial systems, which makes it an ideal candidate for a comprehensive simulation platform. Nuclear power plants place a strong emphasis on safety and reliability constraints, which introduce many scenarios that are also relevant to other industrial control systems. Energy Encyclopedia’s work [27] significantly enhanced the digital twin community by providing basic models for different types of nuclear power plants. Models and map configurations of the Unreal Engine environment are modified to synchronize a physics simulator of nuclear power plants, called Generic Pressurized Water Reactor (GPWR) Simulator [28] developed by Western Services Corporation (WSC). The iFANnpp digital twin model is divided into four main zones: the reactor building, the turbine hall building, the water source and cooling towers, and the power area.
@@
Modeled within the reactor building is the reactor pressure vessel and primary loop, including two steam generators, two reactor coolant pumps, and a pressurizer for the primary loop. It also contains industrial components including a fuel loading machine, polar crane, and spent fuel cooling pool. The core contains the reactor fuel rods; inside the core is a mock-up fuel assembly with a 9$\times$9 array of fuel and control rods, which release heat through reactions or moderate reactions, respectively. The GPWR simulator uses a 17$\times$17 fuel assembly array, but the quarter-sized 9$\times$9 array is used in the model because it is simpler to model and less resource-intensive to render during simulations. The core is held by the reactor vessel body, which has inlets for cold water to flow from the reactor coolant pumps and outlets for heated water to flow to the steam generators. Water flows from the reactor coolant pumps through the reactor core, where it is heated up and passes into the steam generator in order to heat up the secondary loop. It then flows to the reactor coolant pump, cools down, and returns to the core, completing the primary loop. The pressurizer is filled with water, which is heated to the boiling point, and maintains pressure in the primary circuit. Fuel loading machine is an equipment to handle the fuel assemblies for loading and unloading operations. The spent fuel system is a special area to store spent nuclear fuel assemblies that have been removed from the reactor core. The polar crane is used to lift and handle major components during various phases of a nuclear power plant’s life cycle, including operation and decommissioning. Lastly, contained in the reactor building are two secondary loop circulation pumps, which pump water in the secondary loop through the steam generators to be heated up by the primary loop, then back to the turbine hall to spin the turbines.
@@
The turbine hall contains the power-generating components of the nuclear power plant to match those of the GPWR simulator: three low-pressure and one high-pressure turbine, an electric generator, and three condensers. The turbines are rotatory devices spun by steam in the secondary loop, effectively converting heat energy into mechanical energy. These turbines rotate inside the electric generator, which converts mechanical energy into electrical energy. As the steam flows through the turbines and cools, it is caught and condenses into water in the condensers, then is released in an external water source or cooled and recycled in the cooling towers. The GPWR simulator models water being released to an external water source and replaced by cooler water upstream of the source. The turbine hall also contains the moisture separator re-heaters, which assist with the condensation process.
@@
The final zone is the power area, where all of the electrical power systems are located. This contains an assemblage of high-voltage power lines that will carry electricity to population centers or recycle power back into the plant.
@@ IV Digital Twin Features
The platform provides user-friendly features, enabling researchers to obtain accurate and detailed results, while easily accessing ongoing processes and outcomes in a single comprehensive view. These features include the Python-Unreal bridge, bidirectional real-time mirroring, robot models, virtual reality, robot navigation utilities, real-time top-down perspective, and thermal vision. Each feature is further explained in subsequent sections.
@@ IV-A Python-Unreal Bridge
While the Python-Unreal bridge provides a high-level form of communication with the Unreal environment, Python is also a popular language used for machine learning. By integrating the ability to communicate with the Unreal Engine, iFANnpp opens up the potential for conducting various experiments while providing access to the entire Python machine-learning ecosystem.
@@ IV-B Bidirectional Real-Time Mirroring
GPWR is a high-fidelity PWR simulator that allows for real-time simulation of complex physical processes in nuclear power plants. In addition, the simulator incorporates physics-based models of fluid, gases, heat transfer, and other phenomena in nuclear power plants. Integrating the Unreal Engine project with the GPWR simulator enhances the framework by giving it a unique characteristic of a Digital Twin: continuous mirroring of real-time data, behavior, and interactions.
Figure 3: Robot Models
To enable the connection between the GPWR simulation software and iFANnpp environment, a Transmission Control Protocol (TCP) client and server architecture are implemented. The server part is a bridge between the GPWR server and the bi-directional pipe client. The client that runs within the Unreal Engine simulation makes requests to the server, which then relays commands to the GPWR simulation software. This procedure allows for reading and writing GPWR variables such as sensor pressure, water level, material weights, and liquid temperature. This connection is also used to update the values of sensors within the GPWR environment. The client polls the server frequently for variable values, allowing for the corresponding values within the Unreal Engine simulation environment to synchronize with the GPWR variables as they change throughout the simulation. As the TCP client code is generic, the PWR simulator could be replaced by a physical PWR system, using the same TCP client. Simulations can be run in real-time to make informed decisions about potential robot actions and the best decision can be confirmed in simulation before deploying it to the real-world. Since the iFANnpp platform synchronizes the state of its connected system with the simulation environment, decisions made in simulation will be based on recent and relevant information.
@@ IV-C Robot Models
A diverse range of robot models is provided within the digital platform to optimize operational capabilities across various tasks in nuclear power plants. Due to the different demands of tasks in such a highly complex and potentially hazardous environment, a variety of mobile systems are required. In iFANnpp, four distinct robot models with varying locomotion capabilities have been imported and developed within the digital twin environment: TurtleBot [robotis_2022] (WheelRobot), Atlas [30] (Bipedal Robot) , Champ [31] (Quadruped Robot) , and Quadcopter [32] (Aerial Robot).
Figure 4: Navigation Utility
Each of these robot types is tailored to specific functions. TurtleBot is designed for tasks requiring high mobility on smooth, flat surfaces and can cover large distances, making it ideal for routine tasks that require frequent movement across different plant sections on smooth surfaces. Atlas mimics human-like locomotion and manipulator controls, allowing it to navigate complex environments such as obstacle areas, stairs, and uneven surfaces, and interact with human equipment. The quadruped robot is constructed for stability and versatility. The four-legged locomotion allows it to maintain balance on uneven terrain and provides agility and durability at the same time. Lastly, Quadcopter can inspect areas that are difficult to reach at altitudes. It can also be used to quickly survey or monitor large areas.
Use of robots within a digital twin can provide multiple potential benefits to the nuclear industry. First, robot models can be simulated and tested in design-specific tasks. The integration of robots with a digital twin allows for the customization and calibration of the robot models based on the specific needs of a given task. Second, developing and testing robot designs in a digital twin can be reduce development time and cost. Engineers can experiment with various robotic designs, features, and workflows in a cost-efficient and risk-free environment. Quick modifications can be made to the robots’ locomotion, sensors, and operational algorithms without the need for expensive prototypes. Lastly, the robots can be tested in potential hazardous areas without direct deployment. Nuclear power plants are highly sensitive environments where safety is critical. By deploying a variety of robots for specific tasks within in a digital twin, the risk associated with such environment can be predicted and eventually avoided.
@@ IV-D Virtual Reality
The incorporation of virtual reality into the platform offers several benefits, especially in the context of nuclear power plant simulation and training. First, VR enables users to engage in hands-on interaction within a fully immersive 3D environment, which makes complex systems or components more straightforward to understand. This immersive experience accelerates learning, especially when requiring extensive real-world setup or lengthy training. Second, VR minimizes the need to consider many variables associated with real-world tests, such as physical prototypes and on-site visits. Traditionally, system testing and training requires constructing expensive physical models or on-site presence. With VR, users can simulate real-world operations with high accuracy and reduce the costs associated with physical development. Furthermore, changes or optimizations in systems or operations can be tested without risk of damaging actual systems. Lastly, VR offers the monitoring of an entire nuclear power plant vividly, allowing trainers to track data and system status in real-time data with clarity and precision.

### v3

@@ Abstract
Robotics has gained attention in the nuclear industry due to its precision and ability to automate tasks. However, there is a critical need for advanced simulation and control methods to predict robot behavior and optimize plant performance, motivating the use of digital twins. Most existing digital twins do not offer a total design of a nuclear power plant. Moreover, they are designed for specific algorithms or tasks, making them unsuitable for broader research applications. In response, this work proposes a comprehensive nuclear power plant digital twin designed to improve real-time monitoring, operational efficiency, and predictive maintenance. A full nuclear power plant is modeled in Unreal Engine 5 and integrated with a high-fidelity Generic Pressurized Water Reactor Simulator to create a realistic model of a nuclear power plant and a real-time updated virtual environment. The virtual environment provides various features for researchers to easily test custom robot algorithms and frameworks.
@@
Digital Twin, Robotics, Nuclear Power Plant, Artificial Intelligence
@@ I Introduction
Robots have emerged as potential transformative tools across many industrial systems including nuclear power plants, driven by their incredible precision [1, 2, 3], enhanced productivity [4, 5, 6], diversity in controls [7, 8, 9], and range of locomotion methods [10, 11, 12]. However, errors in robot control software can result in severe hardware damage or even human injury. In addition, real-world robot development and testing can be slow and expensive. To help mitigate these issues, digital twin platforms can be relied on for safe, rapid, and cost-efficient prototyping.
A digital twin is a solution to develop a realistic, virtual model of a system with continuous updates of its real-time data and behavior. It enables real-time monitoring and physics simulation, allowing operators to predict system behaviors. The increase in predictability improves nuclear power plant reliability by reducing plant failures. Additionally, a digital twin provides a platform for testing research ideas without accessibility constraints. Access to nuclear power plants is highly restricted due to their potential hazards and confidentiality. A digital twin is perfect for this environment since it presents the simulation of the nuclear power plants and produces data that can be further used for multiple functions. Additional follow-up research can be conducted, including simulation of various operations in nuclear power plants, optimization of the operations, prediction of outcomes using robots, and real-time monitoring of nuclear power plant component behaviors.
Figure 1: iFANnpp architecture and features
This paper introduces a nuclear power plant digital twin for robots and the development of artificial intelligence (AI), called iFANnpp. The digital twin is capable of supporting the development of advanced technologies such as machine learning (ML), reinforcement learning (RL) and autonomous robot controls. Figure 1 illustrates the overview framework of iFANnpp, which includes the core digital twin features such as a bidirectional communication pipeline using a high-fidelity physics simulator for accurate nuclear power plant simulation and a Python-Unreal bridge for a user-friendly ML/RL testbed. This platform uses Unreal Engine 5 (UE5), a powerful 3D modeling and simulation software. It offers several key features:
@@
The remainder of the paper is structured as follows: Section II introduces related research on industrial digital twins and associated robots. Section III discusses the modeling of iFANnpp. Section IV describes the various features presented in iFANnpp. Section V demonstrates potential use cases of the digital twin. Section VI presents an analysis on the performance of the digital twin. Section VII concludes the paper.
@@ II-A Industrial Digital Twin
Several studies have explored integrating advanced 3D models with real-time data to create accurate and functional digital twins. Sørensen et al. [14] developed a digital twin using Unreal Engine 5 specifically for wind power analysis, allowing real-time simulation and turbine performance optimization. Garg et al. [15] implemented a virtual reality (VR) feature with Unity 3D to connect a virtual world to a physical FANUC robot, providing remote control and monitoring its operation. Erdei et al. [16] used Unreal Engine 4 to create a cyber-laboratory space for KUKA KR5 work area for training purposes. Metzner et al. [17] created a VR digital twin of a robot in an industrial control system, specifically for simulating programmable logic controllers (PLCs) in Unity. Gonzalez et al. [18] modeled a laboratory room virtually and augmented it with reality capabilities using a mobile device, enhancing virtual interaction more realistically. Vairagade et al. [19] developed an NVIDIA Omniverse-based digital twin for nuclear power plants using the same Generic Pressurized Water Reactor (GPWR) simulator, a prior work that was limited to realistically modeling a single building within the nuclear power plant system. This work models all primary functional components of a pressurized water reactor (PWR) and offers a broader range of features, including bidirectional communication with the underlying simulator and the Python-Unreal bridge.
@@ II-B Robots in Industrial Environments
While the former works focus on the modeling and control of stationary robots to enable fine-grained manipulation and automation for industrial applications, additional works feature mobile robotics. Mobile robotics offer the flexibility of movement, but can also add design complexity regarding power management and algorithmic control [23]. Naveen et al. [24] designed and implemented a solar-powered mobile robot with pick and place capabilities as well as control software for potential industrial automation tasks. Ghodsian et al. [25] provides a survey of industrial mobile robotics, highlighting the importance of their development for Industry 4.0 designs, their flexibility in terms of movement and coordination, as well as the complexities and challenges in their design to ensure valid function and safety. Buerkle et al. [26] discuses a potential industrial robot system architecture using mobile robots, called Industrial Robots as a Service, highlighting potential methods for programming mobile robots such as through explicit task programming or via human-robot interaction. Additionally, the work emphasizes the need for robot adaptability to environmental change and the critical importance of safety, primarily when humans work near autonomous machines.
iFANnpp provides a platform for simulating stationary and mobile robots specifically for nuclear power plant applications. The transition to deploying mobile robots for industrial applications presents many challenges, including the need for more complex algorithm development and a safe environment. Using iFANnpp, researchers can prototype mobile robot planning and control methods, confirming design and execution validity before deploying them to the real world.
@@ III Digital Twin Modeling
iFANnpp includes 3D models for the entire nuclear power plant, shown in Figure 2. A nuclear power plant is a complex industrial system, which makes it an ideal candidate for a comprehensive simulation platform. Nuclear power plants place a strong emphasis on safety and reliability constraints, which introduce many scenarios that are also relevant to other industrial control systems. This work utilized basic models for an PWR from Encyclopedia [27] and modified it and expanded it to synchronize a high fidelity full scope GPWR Simulator [28] developed by Western Services Corporation (WSC). The iFANnpp digital twin model is divided into four main zones: the reactor building, the turbine hall building, the water source and cooling towers, and the power area.
@@
Modeled within the reactor building is the reactor pressure vessel and primary loop, including two steam generators, two reactor coolant pumps, and a pressurizer for the primary loop. It also contains industrial components including a fuel loading machine, polar crane, and a spent fuel cooling pool. The core contains the reactor fuel rods; inside the core is a mock-up fuel assembly with a $\times×9$ array of fuel and control rods, which release heat through reactions or moderate reactions, respectively. The GPWR simulator uses a $\times×17$ fuel assembly array, but the quarter-sized $\times×9$ array is used in the model because it is simpler to model and less resource-intensive to render during simulations. The core is held by the reactor vessel, which has inlets for cold coolant to pumped by reactor coolant pumps and outlets that connects to hot leg and then to the steam generators. Coolant is pumped by the reactor coolant pumps into the reactor core, where it is heated up and enters into the steam generator to heat up the secondary loop. The pressurizer maintains pressure in the primary loop. The fuel loading machine handles the fuel assemblies for the loading and unloading operations. The spent fuel system is a special area for storing spent nuclear fuel assemblies that have been removed from the reactor core. The polar crane is used to lift and handle major components during various phases of a nuclear power plant’s life cycle, including operation and decommissioning.
@@
The turbine hall contains the power-generating components of the nuclear power plant to match those of the GPWR simulator: three low-pressure and one high-pressure turbines, an electric generator, and three condensers. The turbines are rotatory devices spun by steam in the secondary loop, effectively converting heat energy into mechanical energy. These turbines drives the electric generator, which converts mechanical energy into electrical energy. The moisture separator re-heaters are also modeled to improve the steam quality before enters into the low pressure turbine. As the steam flows through the turbines and becomes exhaust steam, it is then condensed into feedwater in the condensers. Feedwater then pumped through different stage of reheaters and then enters into the steam generator secondary side.
@@
The final zone is the power area, where all of the electrical grid transmission systems are located. This contains an assemblage of high-voltage power lines and transformers that will carry electricity to the power grid, which ultimately supplies power to cities.
@@ IV Digital Twin Features
The platform provides user-friendly features, enabling researchers to obtain accurate and detailed results, while easily accessing ongoing processes and outcomes in a single comprehensive view. These features include the Python-Unreal bridge, bidirectional real-time mirroring, robot models, virtual reality, robot navigation utilities, real-time top-down perspective, and thermal vision. Each feature is further explained in the subsequent sections.
@@ IV-A Python-Unreal Bridge
While the Python-Unreal bridge provides a high-level form of communication with the Unreal environment, Python is also a popular language used for AI and ML. By integrating the ability to communicate with the Unreal Engine, iFANnpp opens up the potential for conducting various experiments while providing access to the entire Python AI/ML ecosystem.
@@ IV-B Bidirectional Real-Time Mirroring
GPWR is a high-fidelity pressurized water reactor simulator that allows for real-time simulation of complex physical processes in nuclear power plants. In addition, the simulator incorporates physics-based models of fluid, gases, heat transfer, and other phenomena in nuclear power plants. Integrating the Unreal Engine project with the GPWR simulator enhances the framework by giving it a unique characteristic of a digital twin: continuous mirroring of real-time data, behavior, and interactions.
To enable the connection between the GPWR simulation software and iFANnpp environment, a Transmission Control Protocol (TCP) client and server architecture are implemented. The server part is a bridge between the GPWR server and the bidirectional pipe client. The client that runs within the Unreal Engine simulation makes requests to the server, which then relays commands to the GPWR simulation software. This procedure allows for reading and writing GPWR variables such as sensor pressure, water level, material weights, and liquid temperature. This connection is also used to update the sensor values within the GPWR environment. The client polls the server frequently for variable values, allowing for the corresponding values within the Unreal Engine simulation environment to synchronize with the GPWR variables as they change throughout the simulation. As the TCP client code is generic, the simulator could be replaced by a physical PWR system, using the same TCP client. Simulations can be run in real-time to make informed decisions about potential robot actions, and the best decision can be confirmed in simulation before deploying it to the real world. Since the iFANnpp platform synchronizes the state of its connected system with the simulation environment, the decisions made in simulation will be based on recent and relevant information.
@@ IV-C Robot Models
Figure 3: Robot models (A) wheeled (B) aerial (C) humanoid (D) quadruped
A diverse range of robot models is provided within the digital platform to optimize operational capabilities across various tasks in nuclear power plants. Due to the different demands of tasks in such a highly complex and potentially hazardous environment, a variety of mobile systems are required. In iFANnpp, four distinct robot models with varying locomotion capabilities have been imported and developed within the digital twin environment: TurtleBot [29] (WheelRobot), Atlas [30] (Bipedal Robot) , Champ [31] (Quadruped Robot) , and Quadcopter [32] (Aerial Robot).
Each of these robot types is tailored to specific functions. TurtleBot is used for tasks requiring high mobility on smooth, flat surfaces. It is ideal for routine tasks that require frequent movement across different plant sections on smooth surfaces. Atlas mimics human-like manipulator controls, allowing it to interact with human equipments. Champ is utilized in a situation where it requires stability in locomotion. The four-legged locomotion allows it to maintain balance on uneven terrain and provides agility and durability at the same time. Lastly, a quadcopter can inspect areas that are difficult to reach at altitudes. It can also be used to quickly survey or monitor large areas.
Use of robots within a digital twin can provide multiple potential benefits to the nuclear industry. First, robot models can be simulated and tested in design-specific tasks. The integration of robots with a digital twin allows for the customization and calibration of the robot models based on the specific needs of a given task. Second, developing and testing robot designs in a digital twin can reduce development time and cost. Researchers can experiment with various robot types, features, and workflows in a risk-free environment. Quick modifications can be made to the robots’ locomotion, sensors, and operational algorithms without the need for expensive prototypes. Lastly, robots can be tested in potential hazardous areas without direct deployment. Nuclear power plants are highly sensitive environments where safety is critical. By deploying a variety of robots for specific tasks in a digital twin, the risk associated with such environment can be predicted and avoided.
@@ IV-D Virtual Reality
The integration of VR into the digital twin offers several benefits, especially in the context of nuclear power plant simulation and training. First, VR enables users to engage in hands-on interaction within a fully immersive environment. This makes complex systems or components more straightforward to understand. This immersive experience accelerates training, especially when requiring extensive real-world setup. Second, VR minimizes the need to consider many variables associated with real-world tests, such as physical prototypes and on-site visits. Traditionally, system testing and training requires constructing expensive physical models or on-site presence. With VR, users can simulate real-world operations with high accuracy and reduce the costs associated with physical development. Lastly, changes or optimizations in systems or operations can be tested without risk of damaging actual systems.